# N-GLARE: An Non-Generative Latent Representation-Efficient LLM Safety Evaluator

### v1

@@ 4.2 Benign Manifold and Geometric Turning Angle
In practice, since $n^{(G)}(h)$ is collinear with $r^{(G)}(h)$, we compute the numerator as $\tau_{t}^{(A,G)\top}r^{(G)}\bigl(h_{t}^{(A,G)}\bigr)$ and normalize accordingly. Intuitively, $\theta_{t}^{(A,G)}$ measures how strongly the trajectory at $t$ turns “away” from the benign manifold. Smaller angles (particularly $\theta_{t}^{(A,G)}<\pi/2$) indicate movement into regions further from benign behavior.
Algorithm 1 JSS Core Computation for Model $M$
@@
20:    if $s_{i}\leq s_{0}$ then
21:     $\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})\mathrel{+}=\mathrm{JS}_{i}^{(G)}(A,B^{\prime})$
22:    end if
@@
24:  end for
25:  $\mathrm{JSS}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B^{\prime})$
26:  $\mathrm{JSS}_{\text{early}}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})$
27:end for
28:return $\mathrm{JSS}(J,B)$, $\mathrm{JSS}_{\text{early}}(J,B)$, $\mathrm{JSS}(J,R)$, $\mathrm{JSS}_{\text{early}}(J,R)$, $\mathrm{JSS}(P,B)$
4.3 Probing Conditions, Slicing, and JSS Metrics
We apply the above constructions to three probing conditions that are designed to be isomorphic along the progress axis $s$:
  • Benign ($B$): natural or conventional task distributions, used to fit the benign manifold and serve as a reference.
  • Jailbreak ($J$): adversarial sequences designed to elicit harmful behavior.
  • Ideal refusal ($R$): sequences identical to $J$ in user content, but with the assistant response replaced by a fixed, policy-compliant refusal template.
  • Plain Query ($P$): sequences that repeatedly present the raw jailbreak true intention to the model, capturing the model’s unconstrained internal response trajectory under direct exposure to harmful prompts.
For each group $G$ and condition $A$, we thus obtain a set of angle–progress pairs (i.e. Angular-Probabilistic Trajectories, APT)
$\Theta^{(A,G)}=\{(\theta_{t}^{(A,G)},s_{t})\}.$ (19)
To study how separability evolves along the interaction, we discretize the standardized axis $s\in[0,1]$ into $I$ slices (bins) $\{s_{i}\}_{i=1}^{I}$. In slice $s_{i}$, we collect all angles whose progress falls into that interval:
$\Theta_{A}^{(G)}(s_{i})=\{\theta_{t}^{(A,G)}:s_{t}\in s_{i}\}.$ (20)
We also define an early-stage threshold $s_{0}\in(0,1)$ (default $s_{0}=0.4$) and use it to construct “early-ss” metrics that focus on whether the model exhibits safe behavior in the initial portion of the trajectory.
Within each slice $s_{i}$ and group $G$, we estimate the empirical distributions of the angle samples for two conditions $A$ and $B$ (e.g., $J$ vs. $B$, or $J$ vs. $R$), denoted by $\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))$ and $\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))$, meaninig the distribution of them. We then compute the Jensen–Shannon divergence:
$\begin{split}\mathrm{JS}_{i}^{(G)}(A,B)=\mathrm{JS}\bigl(\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))\,\big\|\,\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))\bigr)\end{split}$ (21)
To obtain a model-level separability measure, we aggregate slice-level divergences. For each group $G$,
$\mathrm{JSS}^{(G)}(A,B)=\sum_{i=1}^{I}\,\mathrm{JS}_{i}^{(G)}(A,B),$ (22)
and average across layer groups to obtain the all-s separability:
$\mathrm{JSS}(A,B)=\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B).$ (23)
Restricting to slices with $i\leq s_{0}$ yields the early-s separability:
$\mathrm{JSS}_{\text{early}}(A,B)=\frac{1}{|G|}\sum_{G}\sum_{i\leq s_{0}}\mathrm{JS}_{i}^{(G)}(A,B).$ (24)
Algorithm 1 summarizes the core computation of JSS metrics for a single model $M$. For clarity, the pseudocode illustrates a single run on the full dataset. In our experiments, we wrap this procedure in a bootstrap loop to obtain empirical confidence intervals and perform non-parametric statistical tests across models.
(a) Lower Group
(b) Middle Group
(c) Upper Group
(d) Lower Group
(e) Middle Group
(f) Upper Group
Figure 2: APT trajectories collected from different layer groups. Subfigures (a)(b)(c) correspond to safety-aligned models, while (d)(e)(f) are from jailbreak-tuned models. Red curves represent Plain Query trajectories, yellow denote Benign trajectories, blue indicate Jailbreak trajectories, and green correspond to Ideal Refusal trajectories. Safety-aligned LLMs exhibit well-separated and low-variance APT trajectories, whereas jailbreak-tuned LLMs show weakened separability and substantially higher variance across all query types, indicating degraded safety boundaries.
5 Experiments and Analysis
This section begins by presenting the intuitive safety characteristics derived from latent representation trajectories, illustrating how model representations evolve under different dialogue families. Subsequently, from a dynamical perspective, we systematically examine the correspondence between these latent features and the model’s surface-level safety behaviors, thereby validating the safety representation hypothesis proposed earlier. On this basis, we employ quantified safety feature metrics to evaluate various aligned models and compare their rankings with those obtained from systematic traditional red-teaming evaluations, demonstrating a high degree of consistency in safety scales as well as a significant advantage in evaluation cost. Finally, we investigate the statistical consistency of the ranking results across different datasets and sample sizes, analyzing the stability of this evaluation framework under changes in data distribution and sampling conditions.
5.1 Trajectory Evolution Dynamics
This subsection empirically establishes one core observation of our evaluation framework:
Observation 1.
A model’s safety state is not only a behavior-based static attribute, but rather a dynamic process that can be continuously tracked within the latent representation space.
Specifically, in Figure 2, we compare the APT evolutionary trajectories of the PlainQuery, Baseline, Jailbreak, and Ideal Refusal dialogue families. Baseline maintains a stable, flat trajectory, representing the model’s consistent generation direction during natural dialogue. PlainQuery rises sharply across all layers, rapidly diverging from the baseline, indicating that the model promptly shifts its generation direction upon detecting harmful intent, forming a distinct refusal response. Jailbreak trajectories gradually diverge upward from the baseline, exhibiting a notable downward drift in the higher layers, suggesting that adversarial semantics accumulate in upper-layer representations, progressively eroding the refusal mechanism. Although the Ideal-Refusal trajectory initially follows the Jailbreak path, a clear bifurcation emerges, indicating that the model encodes a latent distinction between accepting and refusing harmful instructions. Such separation suggests that the model possesses an intrinsic cognitive bias that differentiates unsafe compliance from defensive rejection within its internal representation space.
@@ 5.1 Trajectory Evolution Dynamics
Figure 3: Multi-perspective cross-model safety ranking consistency between N-GLARE and traditional red-teaming. For each model family ((a–f; see Appendix Table II for model abbreviations on scatters)), the left scatter plot compares the relative ranks induced by our JSS-based mixed proxy (X-axis) and by Hybrid red-teaming scores (Y-axis); points lying near the diagonal blue band indicate close agreement between the two rankings. The right panel in each family shows the corresponding raw JSS-proxy (blue) and Hybrid (red) scores over all subset models, with shaded regions denoting bootstrap variability, again yielding almost identical within-family orderings. Note that ”Hybrid” is the overall average across score all tested red-teaming tasks, with detailed task-wise scores provided in Appendix Table LABEL:tab:jb_safety_multi.
5.2 Synchronous Evolution of Jailbreak Progression, Refusal Tendency, and JSS Dynamics
After obtaining the intuitive empirical phenomena, we focus on verifying two key observations:
Observation 2.
Divergence between Jailbreak and Baseline trajectories as an indicator of the model’s discriminative capability to detect jailbreak attacks process shifting.
A separation between the jailbreak attack trajectory and the baseline behavior trajectory in the representation space signifies a clear process shift in the model’s internal dynamics from normal generation to adversarial alignment. This divergence reflects the model’s discriminative capability to detect and respond to emerging jailbreak cues, serving as a direct and quantifiable indicator of its latent safety awareness.
Observation 3.
Distance between Jailbreak and Ideal Refusal trajectories as an indicator of the model’s intrinsic refusal tendency.
The JS distance between Jailbreak and Ideal Refusal trajectories directly reflects the model’s ability to distinguish and resist jailbreak behaviors. An increase in this distance indicates that the model has a stronger latent discrimination between “accepting” and “blocking” jailbreak instructions, thus exhibiting an enhanced intrinsic refusal tendency, corresponding to a higher level of safety awareness.
5.2.1 Phase-wise Dynamics of Jailbreak Trajectories and Latent Safety Awareness
For observation 2, if a model lacked latent safety awareness, JB activations would remain near the Baseline manifold, yielding no meaningful temporal segmentation. To test whether phase structure exists and whether it is safety-dependent, we use a lightweight neural classifier as an operational probe: if representations at different jailbreak stages encode distinct information, the classifier should reliably separate them. Aligned Llama-family models indeed show strong separability: 0.86–0.97 (start vs. mid), 0.91–0.96 (start vs. end), 0.85–0.86 (mid vs. end). However, after ablating safety alignment (manually jailbroken variants), stage classification collapses to $\sim 0.33$ for all pairs. This contrast provides two conclusions: (1) the phase structure is statistically reliable, because it can be consistently decoded by a simple classifier; (2) it is safety-dependent, because the structure disappears once alignment is destroyed. Thus, aligned models internally encode a latent variable corresponding to the stage of the jailbreak process, and the resulting phase-wise divergence of JB from Baseline offers a direct, graded indicator of latent safety awareness.
5.2.2 Distance between Jailbreak and Ideal Refusal Trajectories and the Model’s Refusal Tendency
For another observation 3, empirical measurements confirm a strong synchrony between Ideal Refusal-Jailbreak separability and model’s refusal tendency. We evaluate the model’s latent refusal inclination strength by using a unified metric of how strongly the model tends to generate reject-related tokens at each stage of the jailbreak process. Specifically, we formalize ANM, which quantifies probability-level concentration on “refusal-cluster” tokens.
$\mathrm{ANM}(x)=\big\langle W_{\mathrm{ref}},\,p(\cdot\mid x)\big\rangle=\sum_{v\in\mathcal{V}}W_{\mathrm{ref}}(v)\,p(v\mid x),$ (25)
where $p(v\mid x)=\mathrm{softmax}(z_{v}(x)/\tau_{\text{pred}})$ is the model’s next-token distribution, and $W_{\mathrm{ref}}$ is a fixed refusal–prototype distribution constructed by:
$W_{\mathrm{ref}}(v)=\frac{\exp(\langle\hat{e}_{v},\hat{c}\rangle/\tau_{\text{ref}})}{\sum_{u\in\mathcal{V}}\exp(\langle\hat{e}_{u},\hat{c}\rangle/\tau_{\text{ref}})},$ (26)
with $\hat{c}$ obtained by averaging normalized embeddings of several refusal-seed phrases. (”sorry”, ”I can’t”, etc.) Thus, ANM quantifies how strongly the model’s predictive distribution aligns with refusal-semantics in latent space.
Across layers, ANM curves exhibit a striking “mid-phase rise and end-phase collapse” pattern that matches the Ideal Refusal-Jailbreak JSS waveform projected into the same 0–1 domain. According to the results, layer-wise correlations between ANM and Ideal Refusal-Jailbreak JSS (r=0.70,0.56,0.54;RMSE=0.130.20r=0.70,0.56,0.54;\ $\text${RMSE}=0.13–0.20)[hastie2009elements] reveal a strong phase-locked coupling between representational separability and refusal intensity. In each layer, the mid-segment elevation ($s\approx 45–65$) corresponds to a surge of refusal semantics, while the late-stage decline ($s>70$) aligns with the erosion of this tendency as jailbreak success emerges.
The consistent alignment between ANM and Ideal Refusal-Jailbreak separability provides a quantitative bridge between latent-space dynamics and explicit refusal behavior, establishing an interpretable and measurable indicator of the model’s latent safety awareness.
5.2.3 Conclusion
Overall, we demonstrate that a model’s safety is manifested not only in its observable output content, but more fundamentally in the organization and evolution of its latent representation trajectories under our method. Therefore, analyzing these trajectory’s separability can equally reveal the core safety mechanisms that traditional output-based evaluations seek to capture.
5.3 Model Differentiation and Safety Characterization
This section presents the core capability of our evaluation framework: its ability to precisely and transferably characterize safety differences across models. In other words, the JSS-based system not only captures the internal representational dynamics of a model but also enables multi-scale and fine-grained quantification of safety variation and underlying regularities across different alignment levels. Within this overarching framework, our analyses proceed along two complementary axes:
  • Inter-Model Dimension (Model-Level Differentiation): examining subtle variations in latent-space separability among instruction-aligned models to reveal the continuous distribution of intermediate safety regimes;
  • Intra-Model Dimension (Weight-Level Evolution): contrasting forward DPO and inverse DPO settings to observe how safety alignment evolves directionally with changes in alignment weight.
Together, these two perspectives form a progressive precision hierarchy: from inter-model to intra-model analysis—demonstrating that the JSS framework consistently reflects alignment strength across granularities, achieving a unified, continuous, and model-invariant scale for safety evaluation.
Figure 5: Layered subset rank consistency with statistically significant (bottom right) agreement across all $\tau$ thresholds.
5.3.1 Cross-Model Validation of JSS-Derived Safety Proxies
For model-level differentiation, we collected a diverse set of models and evaluated their performance under a series of standardized red-teaming safety tests (21 testing protocols with over 7000 test cases). For each model, the results were aggregated into a single averaged safety score (Hybrid Score). Our goal in this experiment is twofold: (1) to derive proxy safety indicators from our JSS-based evaluation framework, and (2) to assess the consistency between the proxy-based ranking and the ranking obtained from conventional red-teaming safety evaluations.
A higher value indicates that the overall separation strength between Jailbreak and Baseline trajectories is significant based on the metric that between PlainQuery and Baseline. This reflects stronger discriminability of jailbreak trajectories.
A higher ratio indicates that the Ideal Refusal-Jialbreak separability remains stable across the entire trajectory. If the separation persists throughout decoding, the minimum approaches the maximum, resulting in a higher ratio. Conversely, if the separation collapses near the end—implying successful jailbreak intrusion and erosion of refusal—this ratio drops sharply, capturing the characteristic “mid-phase emergence–end-phase collapse” dynamics. To obtain a unified scalar JSS-proxy for safety alignment evaluation, we combine the two secondary metrics with equal weighting:
$\mathrm{JSS\text{-}Proxy}=0.5*\mathrm{JB/PB\ Ratio}+0.5*\mathrm{JR\ Min/Max}.$ (29)
We quantify the agreement between our proxy-based safety ranking $r_{\text{jss-proxy}}$ and the ranking obtained from conventional red-teaming evaluations $r_{\text{Hybrid}}$ using Kendall’s $\tau$ correlation coefficient:
$\tau(r_{\text{jss-proxy}},r_{\text{Hybrid}})\in[-1,1].$ (30)
Figure  5 and  3 show that cross a sufficiently large model set, the proxy-based ranking exhibits strong concordance with the red-teaming ground truth. This demonstrates that our JSS-derived proxy indicators can accurately reproduce traditional safety orderings within representative alignment regimes, while requiring far lower evaluation cost. As the model pool expands, $\tau$ shows a mild decrease but remains consistently high, with $p$-values far below conventional significance thresholds. This indicates that even under broader model coverage, the alignment between the proxy-based and red-teaming rankings remains statistically significant. Such behavior illustrates the robust hierarchical consistency of our proxy system, maintaining statistically significant agreement at scale while achieving near-perfect correspondence within focused model subsets.
(a) Inverse DPO Baseline-Jailbreak
(b) Inverse DPO Jailbreak-Plian refusual
(c) Inverse DPO Jialbreak-Refuse existing
(d) Forward DPO Baseline-Jailbreak
(e) Forward DPO Jailbreak-Plian refusual
(f) Forward DPO Jialbreak-Refuse existing
Figure 6: Evolution of JSS and behavioral safety metrics under Forward vs. Inverse DPO. The dynamic evolution of both surface-level safety behavior metrics (solid lines: unsafe rate and refusal rate) and latent proxy metrics (dashed lines: JSS ) as functions of the training global step during the Forward DPO (lower row) and Inverse DPO (upper row) processes. The three columns respectively correspond to separability relations across different dialogue families.
5.3.2 Consistency under Model Weight Variations
In essence, we demonstrate that the geometric safety scale captured by JSS not only reflects static differences among models but also dynamically responds to behavioral transitions induced by weight adjustments. Specifically, we introduce two opposite preference directions: Forward DPO and Inverse DPO, which respectively correspond to the enhancement and attenuation of safety alignment. We employ the preference optimization loss to fine-tune the same base model in two opposite directions:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\,\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\!\big(\beta\,\Delta_{\theta}\big)\Big]$ (31)
$\text{where }\;\Delta_{\theta}=\log p_{\theta}(y^{+}\!\mid\!x)-\log p_{\theta}(y^{-}\!\mid\!x).$ Specifically, we deploy it for two purposes:
Forward DPO. The model is fine-tuned with safety-oriented preferences, where compliant responses are favored over harmful ones ($y^{+}$ = safe, $y^{-}$ = harmful).
Inverse DPO. We reverse the preference direction so that harmful responses are favored over compliant ones ($y^{+}$ = harmful, $y^{-}$ = safe).
To avoid conclusions biased by a single fine-tuning intensity, we systematically generate a structured set of model variants within the LoRA parameter space, fixing the base model and dataset while varying: LoRA target modules and LoRA Rank.
Figure  6 illustrates that from a correlation perspective, the local peaks and troughs of the JSS curves strongly align with fluctuations in the surface safety metrics, including both Unsafe Rate and Refusal Rate (Definition in Appendix Section B). This observation indicates that the degree of separability in latent space evolves in synchrony with the model’s explicit refusal behaviors in output space.
A closer examination reveals a consistent time lag of several training windows between the changes in JSS and the corresponding shifts in surface metrics, with the effect being particularly evident in the early stages of Inverse DPO. Quantitatively, the lag is strongest under Inverse DPO (about 76% of runs exhibit lag). For Forward DPO, in the RefEx–JB family, 74% of trajectories show “JSS leads, refusal lags,” with a median shift of –4.4 steps, whereas in JB–Baseline the effect weakens but persists (64%, median –1.4). For the unsafe rate, both families show near-synchronous behavior (RefEx–JB: 45% lead / 48% lag; JB–Baseline centered around zero with median –0.6).
Together, these results show that declines in latent-space separability reliably precede the corresponding drops in observable safety behaviors—typically by several training windows—establishing JSS as a robust early-warning signal for impending safety degradation during training.
(a) Inverse DPO Pearson
(b) Forward DPO Pearson
(c) Inverse DPO Spearman
(d) Forward DPO Spearmsn
Figure 7: Safety Metrics (UR: Unsafe Rate; $R$: Refusual Rate) and JSS Correlation Heatmap
To complement the dual-line analysis, we compute Pearson and Spearman correlations between latent JSS signals and surface safety metrics across lower, middle, and upper layers (Figure 7). As shown in the heatmaps, the correlation signs are fully consistent across all layer groups and both alignment directions. This confirms that JSS reliably tracks the direction of safety strengthening or weakening, providing a stable latent–surface correspondence.
The goal of this experiment is to quantitatively compare resource consumption of two assessment pipelines under identical task scale and sample size conditions:
  • Ours. Representation-based evaluation using offline attack trajectories. We execute a single forward pass on the subject model $M_{s}$ to extract hidden states; no online generation or additional evaluator inference is required.
  • Baseline (traditional online red-team + llm-based evaluator). We run online red-teaming and safety evaluation using a PyRIT + Ollama setup, involving three model roles:
    • $M_{r}$: red-team attack generator,
    • $M_{s}$: subject model under test,
    • $M_{e}$: evaluator / safety discriminator.
Denote the total number of attack targets by $N$, the average number of turns per session by $\bar{T}$, and the average tokens per step by $\bar{L}$. We formalize token and time costs for both pipelines as follows.
For our method, only a forward pass per turn of offline attack dataset on $M_{s}$ is required and hidden states are cached, then we can get the token cost and runtime cost:
$C_{\text{token}}^{\text{Ours}}(N)\;=\;N\cdot\bar{T}_{\text{offine}},$ (32)
$C_{\text{time}}^{\text{Ours}}\;=\;\mathrm{Time}_{M_{s}^{\text{forward}}}\;+\;\mathrm{Time}_{\text{offline\_JSS}}\;+\;\mathrm{Time}_{\text{benign\_con}},$ (33)
where $\mathrm{Time}_{M_{s}^{\text{forward}}}$ denotes the cost of $C_{\text{token}}^{\text{ours}}(N,J)$, $\mathrm{Time}_{\text{offline\_JSS}}$ denotes the downstream cost for JSS statistics, and $\mathrm{Time}_{\text{benign\_con}}$ denotes time to construct the benign distribution (both $\mathrm{Time}_{\text{benign\_con}}$ and $\mathrm{Time}_{\text{offline\_JSS}}$ are typically much smaller than generation time, so we ignore them in following analysis).
For traditional baseline, each red-team attempt involves a three-way inference flow for $T$ attack turns and $N$ attack targets:
$C_{\text{token}}^{\text{Base}}(N)\;=\;N\cdot\bar{T_{r}}\cdot\big(\bar{L}_{r}\;+\;\bar{L}_{s}\;+\;\bar{L}_{e}\big),$ (34)
where $\bar{L}_{r},\bar{L}_{s},\bar{L}_{e}$ denote the average token lengths for red-team generation, subject-model generation, and evaluator judgment respectively. $T_{r}$ denotes the average turn number of attacks generated by redteaming models.
We run evaluations at multiple red-team intensities $N=\{n_{1},n_{2},\dots\}$ using PyRIT’s redteaming module to generate online attacks, and record produced token counts and elapsed times. For the JSS method, we count a single forward pass per example as one token-equivalent for bookkeeping; the JSS evaluation uses the same original instruction set $N$ corresponding offline multi-turn jailbreak datasets, derived RefEx and PlainR sets. We record token-equivalent costs and elapsed wall-clock time for both pipelines.
We compare the two pipelines along the following axes:
  • Relative share of output tokens: the fraction of total system tokens generated by the subject model $M_{s}$ during evaluation.
  • Total generated-token count: the overall number of tokens produced by all system components during evaluation.
  • Per-token time cost: wall-clock runtime divided by total produced tokens; this metric measures system throughput. Lower per-token time denotes higher efficiency.
Figure 8 summarizes the cost comparison across all metrics, all roles, and all red-teaming depths, the JSS-only pipeline remains dramatically—often orders of magnitude—more efficient than the traditional three-model setup.
Together, these three findings establish that N-GLARE reduces evaluation cost by at least two orders of magnitude while preserving ranking fidelity, effectively removing the generative bottleneck that makes traditional red-teaming expensive, slow, and difficult to scale.

### v2

@@ C.4 Derivation of the Geometric Turning Angle
Since $n^{(G)}(h)$ is collinear with $r^{(G)}(h)$, the angle can equivalently be computed using the residual direction. By construction, $\theta_{t}^{(A,G)}\in[0,\pi]$. Smaller angles indicate that the trajectory is moving more directly away from the benign manifold, corresponding to stronger deviation from normal generation behavior.
This turning angle serves as the local geometric primitive for trajectory-level aggregation in the main text.
Appendix D Cost Model and Efficiency Analysis
This appendix details the cost modeling assumptions, token accounting rules, and runtime decomposition used in the efficiency comparison between the traditional online red-teaming pipeline and our JSS-only evaluation pipeline.
All measurements are conducted under matched task scale and sample size to ensure a fair comparison.
D.1 Evaluation Pipelines and Model Roles
We compare two evaluation pipelines:
JSS-only pipeline (Ours).
Our method operates on offline multi-turn attack trajectories. For each example, we perform a single forward pass on the subject model $M_{s}$ to extract hidden representations. No online generation, attack synthesis, or evaluator inference is required.
Traditional red-teaming pipeline (Baseline).
The baseline follows a standard online red-teaming workflow implemented using PyRIT with Ollama backends, involving three model roles:
  • $M_{r}$: red-team attack generator,
  • $M_{s}$: subject model under evaluation,
  • $M_{e}$: evaluator or safety discriminator.
Each attack session proceeds in multiple turns, with all three models invoked at every turn.
D.2 Token Cost Formulation
Let $N$ denote the number of attack targets, $\bar{T}_{r}$ the average number of attack turns per session, and $\bar{L}_{r}$, $\bar{L}_{s}$, $\bar{L}_{e}$ the average token lengths generated by $M_{r}$, $M_{s}$, and $M_{e}$, respectively.
Baseline.
The total generated-token cost of traditional red-teaming is:
$C_{\text{token}}^{\text{Base}}(N)=N\cdot\bar{T}_{r}\cdot\bigl(\bar{L}_{r}+\bar{L}_{s}+\bar{L}_{e}\bigr).$ (21)
JSS-only pipeline.
Our method requires only a forward pass over an offline multi-turn dataset. For bookkeeping consistency, we count one subject-model forward pass per turn as one token-equivalent unit:
$C_{\text{token}}^{\text{Ours}}(N)=N\cdot\bar{T}_{\text{offline}},$ (22)
where $\bar{T}_{\text{offline}}$ denotes the average number of turns in the offline attack trajectories.
D.3 Runtime Cost Decomposition
Baseline.
The total runtime cost of the traditional pipeline decomposes as:
where each term includes both model inference time and associated system overheads incurred at every attack turn.
JSS-only pipeline.
The runtime cost of our method is:
$C_{\text{time}}^{\text{Ours}}=\mathrm{Time}_{M_{s}^{\text{forward}}}+\mathrm{Time}_{\text{offline\_JSS}}+\mathrm{Time}_{\text{benign\_con}}.$ (24)
Here, $\mathrm{Time}_{M_{s}^{\text{forward}}}$ denotes the time required to perform forward passes on the subject model over the offline dataset. The costs of benign distribution construction and JSS computation are typically negligible compared to model inference and are therefore omitted in efficiency comparisons.
D.4 Measurement Protocol
We evaluate both pipelines at multiple red-teaming intensities $N\in\{n_{1},n_{2},\dots\}$.
For the baseline, online attacks are generated using PyRIT, and all generated tokens from $M_{r}$, $M_{s}$, and $M_{e}$ are counted. Wall-clock time is measured end-to-end for each evaluation run.
For the JSS-only pipeline, the same set of attack instructions is used in the form of offline multi-turn datasets derived from RefEx and PlainR. We record token-equivalent counts and elapsed wall-clock time for the forward pass and offline analysis.
D.5 Evaluation Metrics
We compare the two pipelines using the following efficiency metrics:
  • Total generated-token count: overall system token usage.
  • Subject-model token share: fraction of tokens attributed to the evaluated model $M_{s}$.
  • Per-token time cost: wall-clock time normalized by total generated tokens, serving as a throughput indicator.
Lower per-token time cost indicates higher evaluation efficiency.
D.6 Summary of Efficiency Results
Across all red-teaming intensities and evaluation settings, the JSS-only pipeline consistently reduces both token consumption and runtime by orders of magnitude compared to the traditional online red-teaming pipeline. This reduction arises from eliminating iterative generation and multi-model interaction, thereby removing the primary scalability bottleneck of conventional red-teaming approaches.
Algorithm 1 JSS Core Computation for Model $M$, For clarity, the pseudocode illustrates a single run on the full dataset. In our experiments, we wrap this procedure in a bootstrap loop to obtain empirical confidence intervals and perform non-parametric statistical tests across models.
@@ D.6 Summary of Efficiency Results
20:   end for
21:  end for
22:  $\mathrm{JSS}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B^{\prime})$
@@
24:return $\mathrm{JSS}(J,B)$, $\mathrm{JSS}(J,R)$, $\mathrm{JSS}(P,B)$
Table 4: Temporal lag between JSS changes and Forward DPO surface-level metrics across dialogue families (UR: Unsafe Rate; $R$: Refusal Rate). Lag Prev. denotes the percentage of trajectories in which JSS leads the corresponding surface metric, and Median Shift reports the median temporal offset (in training steps), with negative values indicating that changes in JSS precede changes in the surface metric.
Metric Family Lag Prev. Median Shift
JSS vs. RR RefEx–JB 74 4.4-4.4
JSS vs. RR JB–Baseline 64 1.4-1.4
JSS vs. UR JB–Baseline 53 0.6-0.6
@@
Figure 10: Multi-perspective cross-model safety ranking consistency between N-GLARE and traditional red-teaming. For each model family ((a–f; see Appendix Table 5 for model abbreviations on scatters)), the left scatter plot compares the relative ranks induced by our JSS-based mixed proxy (X-axis) and by Hybrid red-teaming scores (Y-axis); points lying near the diagonal blue band indicate close agreement between the two rankings. The right panel in each family shows the corresponding raw JSS-proxy (blue) and Hybrid (red) scores over all subset models, with shaded regions denoting bootstrap variability, again yielding almost identical within-family orderings. Note that "Hybrid" is the overall average across score all tested red-teaming tasks, with detailed task-wise scores provided in Appendix Table LABEL:tab:jb_safety_benchmark_split_A and Table LABEL:tab:jb_safety_benchmark_split_B.